OpenStudy (anonymous):

need to find the area of trapezoid B with a base of 11 Trapezoid A has an area of 68 with a base of 6 would i use ratio to figure it out

OpenStudy (lgbasallote):

they are similar right?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then yes ratio and proportion

OpenStudy (anonymous):

\[\frac{6}{11}=\frac{68}{B}\]

OpenStudy (anonymous):

124.67

OpenStudy (anonymous):

\[\frac{11*68}{6}=\frac{374}{3}\]